The manager of a restaurant in a large city claims that waiters working in all restaurants in his city earn an average of $150 or more in tips per week. A random sample of 25 waiters selected from restaurants of this city yielded a mean of $139 in tips per week with a standard deviation of $28. Assume that the weekly tips for all waiters in this city have a normal distribution.

 

a. Using the 1% significance level, can you conclude that the manager’s claim is true? Use the p value approach (20 Mks)

 


1
Expert's answer
2021-12-15T16:59:22-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq 150"

"H_1:\mu<150"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=25-1=24" degrees of freedom, and the critical value for a left-tailed test is "t_c = -2.492159."

The rejection region for this left-tailed test is "R = \{t: t < -2.492159\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{139-150}{28/\sqrt{25}}\approx-1.964286"


Using the P-value approach:

The p-value for left-tailed, "df=24" degrees of freedom, "t=-1.964286" is "p = 0.03059," and since "p= 0.03059>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than "150," at the "\alpha = 0.01" significance level.

Therefore, there is enough evidence to claim that the manager’s claim is true, at the "\alpha = 0.01" significance level.